# Haque, Dasgupta, and Rahman

### v1

@@
[…]
Haque, Dasgupta, and Rahman
Grid2Guide: A* Enabled Small Language Model for Indoor Navigation
Md. Wasiul Haque
Department of Civil
   Construction & Environmental Engineering    The University of Alabama
2009 Smart Communities and Innovation Building (SCIB)
   28 Kirkbride Lane
Tuscaloosa
   AL 35487-0288
Email: mhaque16@crimson.ua.edu
Sagar Dasgupta, Ph.D.
Department of Civil
   Construction & Environmental Engineering    The University of Alabama
2009 Smart Communities and Innovation Building (SCIB)
   28 Kirkbride Lane
Tuscaloosa
   AL 35487-0288
Email: sdasgupta@ua.edu
Mizanur Rahman, Ph.D.
Department of Civil
   Construction & Environmental Engineering    The University of Alabama
2007 Smart Communities and Innovation Building (SCIB)
   28 Kirkbride Lane    Tuscaloosa    AL 35487-0288
Email: mizan.rahman@ua.edu
1 Abstract
@@ 2 Introduction
Navigating complex indoor environments can be a time-intensive and disorienting experience. Airports are the most complex public indoor spaces due to their vast physical footprint, multilevel architecture, and constantly changing operational states. Passengers, especially non-native speakers or those with limited mobility, often struggle to interpret signage, locate gates during tightly scheduled layovers, or identify nearby amenities, specifically when real-time changes, such as gate reassignments or delays, occur [1, 2]. Shopping malls, on the other hand, present a different yet equally challenging navigation problem. These commercial hubs frequently consist of multiple floors, interconnected corridors, and a mix of permanent and temporary installations (i.e., seasonal stalls, kiosks), making the spatial layout complex and less predictable. Visitors often rely on static "you are here" maps, which may even be outdated and difficult to interpret, especially for elderly individuals or tourists [3, 4]. Hospitals, meanwhile, represent a particularly high-stakes use case for indoor navigation due to the urgency, emotional stress, and life-critical nature of the visits. These facilities typically have a non-uniform layout shaped by decades of ad hoc expansions, creating labyrinthine interiors with complex structures, ambiguous signage, and inconsistent floor transitions. This is further complicated for those with cognitive, visual, or mobility impairments, who may need additional support to interpret spatial cues or signage [5, 6]. Several studies have shown that navigation issues are not limited to patients; hospital staff also experience disorientation in large healthcare facilities. For instance, nearly one-third of first-time staff members report confusion about the layout of complex hospital campuses, and the need to frequently provide directions can cost staff hundreds of hours annually [7, 8].
@@
Outdoor navigation has been revolutionized by satellite-based technologies, such as GPS. However, these systems perform poorly indoors due to signal attenuation caused by physical obstacles and barriers, such as walls, ceilings, and structures. As a result, indoor navigation remains a largely unsolved problem, particularly in terms of finding a general, scalable, and user-friendly solution. Existing solutions often rely on infrastructure-heavy methods, such as Wi-Fi fingerprinting, Bluetooth beacons, Radio-Frequency IDentification (RFID), or custom sensor arrays. For instance, platforms like IndoorAtlas and Quupa deliver highly accurate positioning by leveraging magnetic fields or Bluetooth Low Energy (BLE) tags; however, these require dedicated deployment and calibration for each building [9, 10]. While somewhat effective, such systems are expensive to maintain, lack flexibility, and are often impractical for wide-scale adaptation in public venues.
@@
In response to the limitations of traditional indoor navigation systems, researchers have increasingly turned their attention to utilizing floor plans, static architectural diagrams that are already available in most public and commercial buildings, as a scalable and cost-effective foundation for navigation solutions. These maps are typically used for safety signage and navigation, but may contain useful spatial information about the internal layout of the floors of a building. However, accurately interpreting these diagrams requires a level of spatial reasoning and contextual understanding, which is not easily handled in real-time by conventional rule-based systems. All the above needs and limitations underscore the need for navigation support that extends beyond conventional maps and static and dynamic signage, offering more intuitive and personalized guidance.
@@
The emergence of Large Language Model (LLM) agents, such as ChatGPT, Claude, and Gemini, may offer a transformative opportunity in this space. These models are not only capable of interpreting natural language but also have evolved to process and reason over images, making them powerful, multi-modal general-purpose problem solvers. By utilizing their ability to generalize across domains, LLMs can potentially act as an interactive indoor navigation assistant. Instead of relying on expensive infrastructures such as beacons or RFID systems, users could simply upload a floor plan and ask the LLM for directions to a specific room or points of interest (POI). However, this solution also has several drawbacks. To process the floormap image directly using LLM is resource-intensive and time-consuming. Image processing is the main bottleneck of using an LLM with floor maps directly. Even though LLMs have vast generalization capabilities, processing a complex image and extracting meaningful information from it remains a resource-intensive task. Each query can require up to 4-5 minutes to be processed, which is not usable for real-time cases using smart hand-held devices. Even with the most advanced reasoning, the LLM available to date, the accuracy is not particularly sufficient [11]. The high computational requirement issue is more acute because of the limited computing resources of smart hand-held devices.
@@
To address these issues, we present an SLM-based approach that offloads spatial reasoning to a lightweight graph-search technique and utilizes SLM solely to generate human-readable instructions from terse navigation instructions. This approach involves creating an obstacle layout of the indoor map, and then a planning search algorithm can be utilized to find a suitable path from the origin to the destination. The contributions of the paper can be summarized as follows:
@@
    All spatial reasoning of an indoor map is offloaded to A* Search on a binarized occupancy grid, guaranteeing accurate optimal routes in each iteration with low computational resources.
@@
    Interactive and human-understandable navigation instruction is generated with the combination of A* Search and an SLM for a real-time application.
@@
    Integration of A* Search with SLM enables efficient execution on standard commodity CPU/GPU hardware. This computationally optimized pipeline ensures broad deployability on handheld devices, making it well-suited for real-time personal navigation applications.
The rest of this paper is organized as follows. The second section reviews the existing challenges and solutions, and the feasibility of those solutions in indoor navigation presented in different literature. The third section presents A* Search and SLM for indoor navigation. The fourth section presents the experimental setup used to evaluate the proposed pipeline, verifying the processing time and accuracy of generating correct navigation instructions in various test cases. The final section summarizes the usability of A* and LLM for generating precise navigation instructions in indoor environments and provides some directions for future research.
@@ 3 Literature Review
Several studies have examined the applicability of generic Location-Based Services (LBS) platforms, such as IndoorAtlas, Anyplace, Quuppa, and Google Maps for indoor navigation in enclosed environments [12, 13, 14]. While these platforms demonstrate usability within designated areas, they require customized deployment for each specific location [15]. This localized setup process is often time-consuming and resource-intensive, limiting scalability across diverse environments. In contrast, many indoor spaces already possess detailed floor plans, offering a promising opportunity to develop more generalized and scalable navigation solutions that leverage existing infrastructure. Such solutions will eradicate the problems caused by the lack of signals in closed environments.
@@
The current advancements in artificial intelligence (AI) have been greatly reflected in the usage of generative AI tools (e.g., ChatGPT). With great general problem-solving and image processing capabilities, LLMs are utilized to provide generalized solutions in different fields. Leveraging LLM for providing proper navigation instructions in indoor setups is still not widely explored by researchers. One recent study was conducted and observed by Coffrini et al. [11] where the authors have taken 2-directional maps of various indoor locations (i.e., shopping malls and airports), preprocessed the maps, and utilized the LLM to generate specific navigation instructions between two different points. Their experimentation has uncovered some weaknesses in the LLM’s ability to predict accurate navigation instructions. The model often provided unnecessary information and instructed a path that is not physically present, even utilizing a refined system prompt. LLMs can solve many complex tasks, but often are seen to stumble in terms of complex reasoning. Such experimentations can be categorized under ’Few-Shot Prompting’ [16]. The system prompt acts as a high-level directive that defines the model’s role, behavior, and expected response style, thereby shaping the interpretation of few-shot examples. Studies have shown that well-crafted system prompts significantly improve the consistency and contextual accuracy of the model’s outputs [17, 18, 19]. However, this approach is not usable in real-world scenarios due to the slow image processing capabilities of the LLMs. Utilizing the advanced reasoning capability, each prompt takes 4-5 minutes to be processed, creating a significant delay and hindering real-time applicability in hand-held devices [11]. The computationally expensive method also introduces the requirement for costly, computationally resource-intensive infrastructure.
@@
Finding an optimized route in an indoor environment can be categorized as a classic path planning from. Efficient graph searching algorithms have been proposed in path planning scenarios in various aspects [20]. In such approaches, an alternative graph-like map of the environment is created to invoke existing graph searching algorithms. Various methods have been proposed to model the environment, such as Voronoi diagrams [21], visibility graphs [22], and grid structures [23]. Conventional shapes are used to represent the environment in the grid structure method, and each grid is consistent, making it suitable for most indoor use cases. The path search algorithm is a method used to find the path between the origin and destination on a map. Based on the approaches, such algorithms can be divided into bionic algorithms, geometric algorithms, and interpolation-based algorithms. Bionic algorithms are random search algorithms simulating natural biological evolution. A group of algorithms, such as the particle swarm optimization (PSO) algorithm [24], the genetic algorithm (GA) [25], and the ant colony (AC) algorithm [26], have been applied to path search processes. Even though these are easy to implement, they often suffer from local optima, making them inappropriate for path planning in complex grid map environments. Interpolation-based algorithms utilize smooth curves, such as the B-spline curve [27] and Dubins curves [28] to fit the entire path. The generated paths are smooth, but they do not fit well to irregular paths; hence, their application is mainly suitable for smoothing local or simple path problems. On the other hand, geometric search algorithms are widely used as path search algorithms. Most common examples include the Dijkstra algorithm [29] and the A* algorithm [30]. The Dijkstra algorithm obtains the shortest path by traversing all the nodes in every iteration, which is computationally expensive. The A* algorithm is based on the Dijkstra algorithm that optimizes by adding a heuristic function, which was first proposed by Hart [31, 32]. The A* algorithm is suitable for real-time usage due to its faster performance, and is therefore widely used in path planning.
@@
In recent years, compact instruction-tuned transformer models have emerged as powerful tools capable of following complex prompts with high fidelity, while maintaining efficiency in terms of computation and memory. This makes them particularly suitable for real-time applications such as indoor navigation, where system responsiveness and resource constraints are critical. Instruction tuning has been shown to substantially improve the ability of language models to generalize across a variety of tasks. Wei et al. demonstrated that instruction-tuned language models substantially outperform their untuned counterparts across a wide range of benchmarks [33]. This finding laid the groundwork for a new generation of efficient language models capable of strong zero-shot and few-shot performance. Google’s FLAN-T5 family delivered state-of-the-art performance with as few as 780 million parameters [34]. This balance between performance and resource efficiency enables use in edge environments such as mobile devices or embedded navigation systems. More recently, Fu et al. evaluated FLAN-T5 (780 M) on real-world meeting summarization and found that it matches or exceeds zero-shot GPT-3.5, despite being an order of magnitude smaller [35]. Other studies have echoed these observations. Frantar et al. introduced GPTQ, a quantization technique that reduces LLM memory footprint without substantial performance loss, enabling models like LLaMA and OPT to run efficiently on consumer hardware [23]. Similarly, Dettmers et al. proposed LLM.int8() as a precision-optimized inference method, allowing larger models to execute on a single GPU with minimal degradation [36]. Moreover, the TinyLlama project has demonstrated that even smaller models, with under 1 billion parameters, can be fine-tuned for high performance in instruction-following tasks while maintaining fast inference speeds and small memory footprints [37]. These models enable seamless deployment on devices with limited compute power, including Raspberry Pi-level hardware or mid-range smartphones, making them ideal for navigation systems that aim to remain lightweight, offline-capable, and privacy-preserving. These indicate that a small, on-device SLM can be utilized to reliably convert the terse text outputs from the A* algorithm to human-understandable navigation instructions, without using any specialized, expensive infrastructure.
@@ 4 Method
Finding a route between two points on a closed circuit with walkable free paths and obstacles is a path planning problem. The initial objective of such is to generate an obstacle-aware map from the provided floorplan image [38]. To achieve this, we convert the image into a binary occupancy grid, classifying each pixel as either walkable or blocked using standard image thresholding and polygon-masking techniques. This preprocessing step yields a discrete 2D grid of free and occupied cells, effectively capturing the navigable space. Once the occupancy grid is generated, the A* search algorithm, augmented with 8-way (diagonal) movement, is employed to compute the shortest path between the specified start and goal coordinates. By offloading geometric reasoning to the A* algorithm, the pipeline eliminates the need for a machine learning model to directly process the image or infer spatial relationships. After obtaining the optimal cell-by-cell route, we apply a lightweight run-length encoding (RLE) scheme to compress repetitive movements and diagonal zig-zag patterns into a concise sequence of directional commands. This compression reduces redundancy and simplifies the input to the final stage.

### v2

# Grid2Guide: A* Enabled Small Language Model for Indoor Navigation

@@
Abstract
@@ 1 Introduction
Navigating complex indoor environments can be a time-intensive and disorienting experience. Unlike outdoor spaces, where landmarks, visibility, and satellite-based tools support orientation, indoor facilities often confine movement to corridors, stairways, or elevators, restricting visibility and increasing reliance on signage and prior familiarity. Among these, airports stand out as some of the most complex public indoor spaces due to their vast physical footprint, multilevel architecture, and constantly changing operational states. In addition to finding gates, passengers must adapt to tightly scheduled layovers, re-routed security checkpoints, and shifting boarding procedures. For travelers with language barriers, cognitive limitations, or mobility impairments, even simple tasks such as locating restrooms or baggage claims can become daunting, especially when compounded by real-time disruptions like gate reassignments or delays Bosch and Gharaveis, (2017); Guerreiro et al., (2019). Shopping malls, on the other hand, present a different yet equally challenging navigation problem. These commercial hubs frequently consist of multiple floors, interconnected corridors, and a mix of permanent and temporary installations (i.e., seasonal stalls, kiosks), making the spatial layout complex and less predictable. Visitors often rely on static ”you are here” maps, which may even be outdated and difficult to interpret, especially for elderly individuals or tourists Belir and Onder, (2013); Jeong et al., (2018). Hospitals, meanwhile, represent a particularly high-stakes use case for indoor navigation due to the urgency, emotional stress, and life-critical nature of the visits. These facilities typically have a non-uniform layout shaped by decades of ad hoc expansions, creating labyrinthine interiors with complex structures, ambiguous signage, and inconsistent floor transitions. This is further complicated for those with cognitive, visual, or mobility impairments, who may need additional support to interpret spatial cues or signage Iftikhar et al., (2021); Jamshidi et al., (2020). Several studies have shown that navigation issues are not limited to patients; hospital staff also experience disorientation in large healthcare facilities. For instance, nearly one-third of first-time staff members report confusion about the layout of complex hospital campuses, and the need to frequently provide directions can cost staff hundreds of hours annually UAB Medicine, (2017); MapsPeople, (2021).
@@
Outdoor navigation has been revolutionized by satellite-based technologies, such as Global Positioning System (GPS). These systems, however, are ineffective indoors because satellite signals experience severe attenuation, blockage, and multipath distortion caused by walls, ceilings, and other structural barriers. As a result, indoor navigation remains a largely unsolved problem, particularly in terms of finding a general, scalable, and user-friendly solution. Existing solutions often rely on infrastructure-intensive methods, such as Wi-Fi fingerprinting, Bluetooth beacons, Radio-Frequency IDentification (RFID), or custom sensor arrays. Other methods leverage on-device sensors through Simultaneous Localization and Mapping (SLAM), which fuses data from cameras, LiDAR, or inertial sensors to construct a map while estimating position. While SLAM avoids external infrastructure, it requires significant computational resources and can struggle in visually repetitive or dynamic environments. Platforms like IndoorAtlas and Quupa deliver highly accurate positioning by leveraging magnetic fields or Bluetooth Low Energy (BLE) tags; however, these require dedicated deployment and calibration for each building Hurtuk et al., (2019); Quuppa, (2019). While somewhat effective, such systems are expensive to maintain, lack flexibility, and are often impractical for wide-scale adaptation in public venues.
@@
In response to the limitations of traditional indoor navigation solutions, researchers have increasingly turned their attention to utilizing floor plans, static architectural diagrams that are already available in most public and commercial buildings, as a scalable and cost-effective foundation for navigation solutions. These maps are typically used for safety signage and navigation, but may contain useful spatial information about the internal layout of the floors of a building. However, accurately interpreting these diagrams requires a level of spatial reasoning and contextual understanding, which is not easily handled in real-time by conventional rule-based systems. All the above needs and limitations underscore the need for navigation support that extends beyond conventional maps and static and dynamic signage, offering more intuitive and personalized guidance.
@@
The emergence of Large Language Model (LLM) agents, such as ChatGPT, Claude, and Gemini, has offered a transformative opportunity in this space. These models are not only capable of interpreting natural language but also have evolved to process and reason over images, making them powerful, multi-modal general-purpose problem solvers. By utilizing their ability to generalize across domains, LLMs can potentially act as an interactive indoor navigation assistant. Instead of relying on expensive infrastructures such as beacons or RFID systems, users could simply upload a floor plan and ask the LLM for directions to a specific room or points of interest (POI). However, this solution also has several drawbacks. To process the floormap image directly using LLM is resource-intensive and time-consuming. Image processing is the main bottleneck of using an LLM with floor maps directly. Even though LLMs have vast generalization capabilities, processing a complex image and extracting meaningful information from it remains a resource-intensive task. Each query can require up to 4-5 minutes to be processed, which is not usable for real-time cases using smart hand-held devices. Even with the most advanced reasoning, the LLM available to date, the accuracy is not particularly sufficient Coffrini et al., (2025). The high computational requirement issue is more acute because of the limited computing resources of smart hand-held devices.
@@
To address these issues, we present a Small Language Model (SLM)-based approach that offloads spatial reasoning to a lightweight graph-search technique and utilizes SLM solely to generate human-readable instructions from terse navigation instructions. This approach involves creating an obstacle layout of the indoor map, and then a planning search algorithm can be utilized to find a suitable path from the origin to the destination. The contributions of the paper can be summarized as follows:
@@
    All spatial reasoning of an indoor map is offloaded to A* Search on a binarized occupancy grid, guaranteeing accurate optimal routes in each iteration with low computational resources. To the best of our knowledge, no prior work has combined classical path planning with LLM/SLM-based indoor navigation, making this a lightweight alternative to resource-intensive LLM reasoning.
@@
    A mechanism for handling multi-layer buildings is incorporated by defining portal nodes (e.g., elevators, escalators, staircases) to enable seamless navigation across multiple floors. To the best of our knowledge, no prior indoor navigation framework has demonstrated lightweight multi-floor path planning combined with natural language instruction generation.
@@
    Interactive and human-understandable navigation instruction is generated with the combination of A* Search and an SLM for a real-time application. Unlike prior systems that rely on static maps or rigid rule-based instructions, this is the first to leverage an SLM for natural language guidance in indoor navigation.
  4. 4.
    Integration of A* Search with SLM enables efficient execution on standard commodity CPU/GPU hardware. This computationally optimized pipeline ensures broad deployability on handheld devices, making it well-suited for real-time personal navigation applications. Existing LLM-based approaches require expensive infrastructure or high-performance computing; to the best of our knowledge, no work has demonstrated such deployability on mobile-class hardware.
The remainder of this paper is organized as follows. Section 2 reviews the existing challenges and current solutions in indoor navigation, highlighting their feasibility and limitations. Section 3 introduces the proposed framework that integrates A* search with a Small Language Model (SLM) for indoor navigation, detailing the underlying components and design. Section 4 describes the experimental setup used to evaluate the proposed pipeline, including test environments, runtime performance, and instruction accuracy. Section 5 discusses the evaluation results, comparing the approach against baseline methods. Finally, Section 6 concludes the paper by summarizing the findings and outlining potential directions for future research.
@@ 2 Literature Review
Several studies have examined the applicability of generic Location-Based Services (LBS) platforms, such as IndoorAtlas, Anyplace, Quuppa, and Google Maps for indoor navigation in enclosed environments Ramani and Tank, (2014); Van der Ham et al., (2016); Wichmann et al., (2024). While these platforms demonstrate usability within designated areas, they require customized deployment for each specific location Furfari et al., (2019). This localized setup process is often time-consuming and resource-intensive, limiting scalability across diverse environments. In contrast, many indoor spaces already possess detailed floor plans, offering a promising opportunity to develop more generalized and scalable navigation solutions that leverage existing infrastructure. Such solutions will eradicate the problems caused by the lack of signals in closed environments.
@@
The current advancements in artificial intelligence (AI) have been greatly reflected in the usage of generative AI tools (e.g., ChatGPT). With great general problem-solving and image processing capabilities, LLMs are utilized to provide generalized solutions in different fields. Leveraging LLM for providing proper navigation instructions in indoor setups is still not widely explored by researchers. One recent study was conducted and observed by Coffrini et al. Coffrini et al., (2025) where the authors have taken 2-directional maps of various indoor locations (i.e., shopping malls and airports), preprocessed the maps, and utilized the LLM to generate specific navigation instructions between two different points. Their experimentation has uncovered some weaknesses in the LLM’s ability to predict accurate navigation instructions. The model often provided unnecessary information and instructed a path that is not physically present, even utilizing a refined system prompt. LLMs can solve many complex tasks, but often are seen to stumble in terms of complex reasoning. Such experimentations can be categorized under ’Few-Shot Prompting’ Sahoo et al., (2024). The system prompt acts as a high-level directive that defines the model’s role, behavior, and expected response style, thereby shaping the interpretation of few-shot examples. Studies have shown that well-crafted system prompts significantly improve the consistency and contextual accuracy of the model’s outputs Reynolds and McDonell, (2021); Ouyang et al., (2022); Wei et al., (2022). However, this approach is not usable in real-world scenarios due to the slow image processing capabilities of the LLMs. Utilizing the advanced reasoning capability, each prompt takes 4-5 minutes to be processed, creating a significant delay and hindering real-time applicability in hand-held devices Coffrini et al., (2025). The computationally expensive method also introduces the requirement for costly, computationally resource-intensive infrastructure.
@@
Finding an optimized route in an indoor environment can be categorized as a classic path planning from. Efficient graph searching algorithms have been proposed in path planning scenarios in various aspects Tang et al., (2021). In such approaches, an alternative graph-like map of the environment is created to invoke existing graph searching algorithms. Various methods have been proposed to model the environment, such as Voronoi diagrams Candeloro et al., (2017), visibility graphs Chou et al., (2019), and grid structures Fransen et al., (2020). Conventional shapes are used to represent the environment in the grid structure method, and each grid is consistent, making it suitable for most indoor use cases. The path search algorithm is a method used to find the path between the origin and destination on a map. Based on the approaches, such algorithms can be divided into bionic algorithms, geometric algorithms, and interpolation-based algorithms. Bionic algorithms are random search algorithms simulating natural biological evolution. A group of algorithms, such as the particle swarm optimization (PSO) algorithm Das et al., (2016), the genetic algorithm (GA) Santiago et al., (2017), and the ant colony (AC) algorithm Xiong et al., (2019), have been applied to path search processes. Even though these are easy to implement, they often suffer from local optima, making them inappropriate for path planning in complex grid map environments. Interpolation-based algorithms utilize smooth curves, such as the B-spline curve Park and Lee, (2007) and Dubins curves Dubins, (1957) to fit the entire path. The generated paths are smooth, but they do not fit well to irregular paths; hence, their application is mainly suitable for smoothing local or simple path problems. On the other hand, geometric search algorithms are widely used as path search algorithms. Most common examples include the Dijkstra algorithm Wang et al., (2011) and the A* algorithm Guruji et al., (2016). The Dijkstra algorithm obtains the shortest path by traversing all the nodes in every iteration, which is computationally expensive. The A* algorithm is based on the Dijkstra algorithm that optimizes by adding a heuristic function, which was first proposed by Hart Hart et al., (1968); Howden, (1968). The A* algorithm is suitable for real-time usage due to its faster performance, and is therefore widely used in path planning.
@@
In recent years, compact instruction-tuned transformer models have emerged as powerful tools capable of following complex prompts with high fidelity, while maintaining efficiency in terms of computation and memory. This makes them particularly suitable for real-time applications such as indoor navigation, where system responsiveness and resource constraints are critical. Instruction tuning has been shown to substantially improve the ability of language models to generalize across a variety of tasks. Wei et al. demonstrated that instruction‑tuned language models substantially outperform their untuned counterparts across a wide range of benchmarks Wei et al., (2021). This finding laid the groundwork for a new generation of efficient language models capable of strong zero-shot and few-shot performance. Google’s FLAN‑T5 family delivered state‑of‑the‑art performance with as few as 780 million parameters Chung et al., (2024). This balance between performance and resource efficiency enables use in edge environments such as mobile devices or embedded navigation systems. More recently, Fu et al. evaluated FLAN‑T5 (780 M) on real‑world meeting summarization and found that it matches or exceeds zero‑shot GPT‑3.5, despite being an order of magnitude smaller Fu et al., (2024). Other studies have echoed these observations. Frantar et al. introduced GPTQ, a quantization technique that reduces LLM memory footprint without substantial performance loss, enabling models like LLaMA and OPT to run efficiently on consumer hardware Fransen et al., (2020). Similarly, Dettmers et al. proposed LLM.int8() as a precision-optimized inference method, allowing larger models to execute on a single GPU with minimal degradation Dettmers et al., (2022). Moreover, the TinyLlama project has demonstrated that even smaller models, with under 1 billion parameters, can be fine-tuned for high performance in instruction-following tasks while maintaining fast inference speeds and small memory footprints Zhang et al., (2024). These models enable seamless deployment on devices with limited compute power, including Raspberry Pi-level hardware or mid-range smartphones, making them ideal for navigation systems that aim to remain lightweight, offline-capable, and privacy-preserving. These indicate that a small, on-device SLM can be utilized to reliably convert the terse text outputs from the A* algorithm to human-understandable navigation instructions, without using any specialized, expensive infrastructure.
@@ 3 Method
Finding a route between two points on a closed circuit with walkable free paths and obstacles is a path planning problem. The initial objective of such is to generate an obstacle-aware map from the provided floorplan image Borges et al., (2019). To achieve this, we convert the image into a binary occupancy grid, classifying each pixel as either walkable or blocked using standard image thresholding and polygon-masking techniques. This preprocessing step yields a discrete 2D grid of free and occupied cells, effectively capturing the navigable space. Once the occupancy grid is generated, the A* search algorithm, augmented with 8-way (diagonal) movement, is employed to compute the shortest path between the specified start and goal coordinates. By offloading geometric reasoning to the A* algorithm, the pipeline eliminates the need for a machine learning model to directly process the image or infer spatial relationships. After obtaining the optimal cell-by-cell route, we apply a lightweight run-length encoding (RLE) scheme to compress repetitive movements and diagonal zig-zag patterns into a concise sequence of directional commands. This compression reduces redundancy and simplifies the input to the final stage.